OpenStudy (anonymous):

What is the solution to the equation -2(1+5a)= -6(a+3) ? a =

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@sleepyjess

OpenStudy (dtan5457):

Distribute and solve for a.

OpenStudy (dtan5457):

Can you attempt to do that for me?

OpenStudy (secret-ninja):

@pielover123 Just do what I taught you before, you move the variable to their own side of the equation and solve it, just like in the last question.

OpenStudy (anonymous):

ok

OpenStudy (secret-ninja):

Its the same thing, just don't over think it. :)

OpenStudy (anonymous):

ok i did it on paper and i got -12

OpenStudy (anonymous):

so would a= -12?

OpenStudy (secret-ninja):

Hmm... not quite... a is 4... Would you mind showing me how you did it?

OpenStudy (anonymous):

ok first what i did is like order of operation

OpenStudy (anonymous):

like i did -2(1+5a)= 6-2=4

OpenStudy (dtan5457):

-2(1+5a)= -6(a+3) When you distribute brackets.. -2-10a=-6a-18

OpenStudy (dtan5457):

How would that be -12?

OpenStudy (anonymous):

then i did (-6a+3)= -3 with would add -3+4= 12

OpenStudy (anonymous):

did i get it right @Secret-Ninja

OpenStudy (secret-ninja):

Lets start from the part where @dtan5457 distributed the brackets. -2-10a=-6a-18 This is where you begin to get the variable (a) by it self.

OpenStudy (anonymous):

ok

OpenStudy (secret-ninja):

-2-10a=-6a-18 Move -2 -10a=-6a-18+2 Move -6a -10a+6a=-18+2

OpenStudy (anonymous):

ok im wirting it on paper

OpenStudy (secret-ninja):

-18+2=-16 and -10a+6a=-4a

OpenStudy (secret-ninja):

So... -4a = -16

OpenStudy (anonymous):

so it is -16

OpenStudy (secret-ninja):

Since -4 is multiplied to a, you divide it to get it to the other side. Tell me, what is -16 divided by -4?

OpenStudy (dtan5457):

a has to be a, not -4a, -a, 4a, nothing, just A.

OpenStudy (anonymous):

ok it is 4 @Secret-Ninja

OpenStudy (dtan5457):

yep. good job everyone.

OpenStudy (anonymous):

so the answer is 4

OpenStudy (anonymous):

thanks guys!!!

OpenStudy (secret-ninja):

Yes, it is 4, good job! :D

OpenStudy (secret-ninja):

No problem. :)